OpenStudy (anonymous):

Solve for x: X- c/2= -c/3

OpenStudy (anonymous):

\[x-\frac{c}{2}=\frac{-c}{3}\] \[x=\frac{-c}{3}+\frac{c}{2}=\frac{c}{6}\]\

OpenStudy (anonymous):

chart is left side = (x-c)/2 or does 2 just divide the c?

OpenStudy (anonymous):

x=-c/3 +c/2 x= (-2c+3c)/6 x=c/6

OpenStudy (anonymous):

2 just divides the 2

OpenStudy (anonymous):

C*

OpenStudy (anonymous):

see above... x=c/6

OpenStudy (anonymous):

Thank you sooo much

OpenStudy (anonymous):

welcome!